OpenStudy (anonymous):

The heights of trees in a certain park are normally distributed with a mean of 30 feet and a standard deviation of 4 feet. What percentage of the trees are between 28 and 36 feet tall? Round the answer to the nearest hundredth.

OpenStudy (anonymous):

P ( 28 < x < 36 ) (28 - 30 ) / 4 = -1/2 ( 36 - 20 )/ 4 = 4 => P ( -.5 < z < 4 )

OpenStudy (anonymous):

Can you look up Z score table now?

OpenStudy (anonymous):

do I look up the Z table for -.5 and 0?

OpenStudy (anonymous):

-.5 and 4

OpenStudy (anonymous):

like -.5 and .04?

OpenStudy (anonymous):

Oops (36 - 30)/ 4 = 1.5

OpenStudy (anonymous):

Just look up Z = .5 and Z = 1.5

OpenStudy (anonymous):

okay so look up the z table for 1.5?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Then P ( 1.5 ) - P ( .5 ) =

OpenStudy (anonymous):

.9332

OpenStudy (anonymous):

do I X that by 30 (the mean)

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

ohh so .9332 is the answer? I meant is there a next step after finding the Z score?

OpenStudy (anonymous):

because the question is asking for a percentage, so I didn't know what to do after i found the z score again.

OpenStudy (anonymous):

z score is .5 and 1.5 You look up from z score to find percent P!

OpenStudy (anonymous):

Then subtract to find the middle!

OpenStudy (anonymous):

so the percent would be 93.32% ?

OpenStudy (anonymous):

Did you subtract P(1.5 ) - P(-.5) ?

OpenStudy (anonymous):

ohhhhhh...I get it now, no. HOld on!

OpenStudy (anonymous):

it equals .2417

OpenStudy (anonymous):

so does that turn into the percent?

OpenStudy (anonymous):

Finally, I can relax to hear "Ohhhh.... " :P

OpenStudy (anonymous):

The table you look it up IS Percentage, Sir!

OpenStudy (anonymous):

haha! not unless I don't pass the test! i have to do it over again. just 3 more days until I graduate so, It's worth it if i have to stay up all night.lol but i'll find someone else to pester with my math questions.

OpenStudy (anonymous):

you mean M'am!

OpenStudy (anonymous):

Ohh, I see ;)

OpenStudy (anonymous):

PASSED! thank you!

OpenStudy (anonymous):

How do you know PASSED?

OpenStudy (anonymous):

Don't forget to round the result to nearest hundreth, though!